# SODAs: Sparse Optimization for the Discovery of Differential and Algebraic Equations

### v1

@@ 3 Methods: Sparse Optimization for Differential-Algebraic Systems
For the identified differential variables, any existing methods for ODE discovery (e.g., SINDy) can be used to find the dynamics equations in the DAE system (Figure 1c). This is the stage where noisy derivatives affect recovery, and there are various approaches to mitigate noise, including weak formulation approaches [16, 17, 18] and smoothing functions [14, 48]. Currently, DaeFinder package supports finding derivatives using finite difference methods and using smoothening filters like splines and Savitzky-Golay filters [48]. The latter is used in our examples. For numerical experiments, we have used LASSO followed by sequential thresholding to solve the sparse optimization. Finally, we obtain a set of differential and algebraic equations of the form (2)–(3) (Figure 1d).
Input: $\mathbf{X}$, $K$, $\Theta$, $\alpha$, sparse_fit, score_function
@@
$k=1$  %set iteration index
condition_number_improved = True   %flag to check if condition number of library has improved.
@@
             sparse_fit$(\Theta_{k-1},\alpha)$   %solve sparse regression problem $\theta_{l}(\mathbf{x})=\sum_{\begin{subarray}{c}j\in\mathcal{I}_{k}\\ j\neq l\end{subarray}}p_{j}^{l}\theta_{j}(\mathbf{x})+\alpha\|\mathbf{p}^{l}\|_{0}$.
             $S^{k}_{l}$ = score_function$(\mathbf{p}^{l},\Theta_{k-1})$
@@
       $m=\arg\max(S^{k}_{l})$  %find the best relationship based on the maximum score.
@@
       $\theta^{(k)}_{\text{high}}=\texttt{select\_term}(g_{k}(\mathbf{x}))$  %select the term inside $g_{k}(\mathbf{x})$ based on the highest complexity ranking.
       $\Gamma_{k}=\texttt{factors}(\theta^{(k)}_{\text{high}},\Theta_{k-1})$   %identify all factors of $\theta^{(k)}$ in $\Theta_{k-1}$.
@@
       condition_number_improved = bool($\text{Cond}(\Theta_{k})-\text{Cond}(\Theta_{k-1})<\epsilon_{k}$)
       $k=k+1$
$\Theta_{*}=\Theta_{k-2}$%final refined library.
Step 3: Dynamical System Finder
@@
$\mathbf{F}(\dot{\mathbf{y}},\ddot{\mathbf{y}},\mathbf{x},\ldots)=0\xleftarrow{}\texttt{discover\_ODEs}(\mathbf{y},\Theta_{*})$   %discovering ODEs using refined library.
@@
Inputs and Parameters: $\mathbf{X}$: Data matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ (time-series data). $K$: Number of previously known algebraic relationships in the library. Default: $K=\texttt{None}$. $\Theta$: Class of candidate library functions. DaeFinder package supports monomials of varying degrees, trigonometric functions,
phase difference terms for power grid, etc.
  • $\alpha$: Regularization parameter for the sparse regression. Default: $\alpha=0.2$. sparse_fit: Function to perform sparse regression, e.g., LASSO, STOLS, etc. Default: LASSO followed by sequential thresholding. score_function: Function to evaluate the quality of sparse fitting. Choices include $\mathcal{R}^{2}$, AIC, BIC, etc. Default: $\mathcal{R}^{2}$. equation_reduce: Function to simplify the best algebraic relationship by taking out common factors in the equation. select_term: Select the highest-ranked term in the reduced equation based on some predefined complexity criteria.
  • Default: for monomial candidate library, degree is used to define rank.
  • factors: Identifies the factors of the selected term inside the library. $\epsilon_{k}$: Tolerance used for stopping the refinement. In practice, this is found using a Pareto front tracking the improvement in $\text{Cond}(\Theta_{k})$. find_dynamic_states: Function to choose dynamic states out of the state space. Various criteria are are discussed in Section 3. discover_ODEs: Function representing any suitable algorithm to discover differential equations. DaeFinder currently supports the
  • strong form of SINDy with smoothening splines [14, 10] and Savitzky-Golay smoothing filters [48].
@@ 4 Results
    We demonstrate the capability of SODAs to discover DAEs and identify reduced coordinate systems through three distinct examples. In Example 1, SODAs rediscovers the underlying DAEs of CRNs with varying complexity, demonstrating the identification of conservation laws and quasi-steady states. Example 2 applies SODAs to pixel data from real-time experiments or animated video of three different cases of nonlinear pendulum: a singular pendulum with negligible damping, a heavily damped pendulum, and a chaotic double pendulum. This example shows SODAs can discover the algebraic constraints and the corresponding reduced coordinate system when such relations are not readily apparent from pixel data. In Example 3,SODAs rediscovers DAEs used in power-grid modeling for three benchmarks (IEEE-4, IEEE-9, and IEEE-39) demonstrating recovery of network topology when the library contains trigonometric functions.
    In Examples 1 and 3, where the time series were generated synthetically, we asses data requirements, robustness to noise, and sensitivity to different types of perturbations are studied. For a simulated time-series states, $\mathbf{X}_{i}$ , Gaussian noise is sampled and added to the time series as
@@
    where $noise\%$ is the percentage of noise, $\sigma_{i}=\text{std}(\mathbf{X}_{i})$, and $\mathcal{N}(0,1)$ is the standard normal random variable. In Example 2, the pixel data was extracted from video footage, which inherently introduced noise into the dataset. In Examples 1 and 2, we explore the impact of polynomial library size on data requirements.
@@ 4.1 Example 1: Chemical Reaction Networks
    We demonstrate the ability of SODAs to accurately recover governing equations from simulated time series of known CRNs, even as system complexity and extrinsic noise increase. In applications in chemistry and biology, the temporal evolution of each chemical species in a CRN can be modeled using mass-action kinetics [49, 50]. This naturally leads to models consisting of polynomial terms of concentration of chemical species. However, timescale separations in the equilibration rates of different species motivate quasi-steady-state approximations [51, 36] which simplify mathematical analysis by reducing the state space dimension while accounting for important interactions and temporal features [52].
    Using SODAs to rediscover the governing equations of CRNs is motivated by unifying the common task of model reduction in chemical and biological networks with model discovery. For known models, a wealth of techniques exist to simplify model representation and reduce complexity. Without loss of information, models can be simplified through conservation analysis, partitioned into sub-networks, or non-dimensionalized. Beyond quasi-steady-state approximations, other reduction techniques include methods that exploit optimization and sensitivity analysis [53]. Although model reduction concepts have previously been associated with neural network-based model discovery [54, 55], SODAs demonstrates these concepts clearly within the framework of regression-based sparse-model selection, leveraging conservation analysis and relevant timescale for interpretable reduction when the model structure is unknown.
    In this example, we use time series generated from three CRNs of varying complexity consisting of one, two, and four enzyme-mediated reactions (Figure 3a). The four-reaction system describes the simplified propanediol utilization pathway native to Salmonella enterica [56, 57]. Details of pathway simulation are included in supplementary Section 1.1. Our data-driven discovery of the pathway at noise levels realistic to those present in metabolite measurement [58, 59] shows the utility of SODAs in detecting the relevant system timescales and offers a guideline on its data requirements and noise robustness.
    Figure 3: Example 1: Application to chemical reaction networks. (a) Nested structure of CRN1 and CRN2 within CRN3 and examples of simulated time-series with $5\%$ noise. Note that only CRN3 allows reversibility of $C\to BE_{2}$ as denoted by the green arrow. (b) Scaling of the data required for recovery of the correct algebraic equations in CRN1 across library degree $p$ and cardinality $|\Theta|$. Non-robust recovery is denoted when fewer than $80\%$ of random seeds succeeded in recovering the correct algebraic relationships. Error bars represent standard error across 10 random noise seeds. (c) Scaling of the data required for recovery of the correct algebraic equations for all CRNs (cardinalities noted) with fixed library degree $p=2$.
@@
    As introduced in the example in Section 2.2, CRNs for the concentrations $\mathbf{x}=[\mathbf{M}]$ of species $\mathbf{M}=[M_{1},M_{2},\dots,M_{N}]^{\top}$ describe reactions between species as
[…]
    $\frac{d\mathbf{x}}{dt}=\mathbf{S}\mathbf{v}(\mathbf{x})$ (22)
[…]
    Algebraic relationships arise in CRNs in two ways. First, the conservation of mass between multiple species considered in the CRN (e.g., enzyme and enzyme-substrate complex) leads to a first-order algebraic relationship (e.g., (17)). Second, a quasi-steady-state approximation resulting from time-scale separation partitions $\mathbf{x}$ into "fast species" $\mathbf{y}$ and "slow species" $\mathbf{z}$, so $\mathbf{x}=[\mathbf{y},\mathbf{z}]^{\top}$ with approximation $\frac{d\mathbf{y}}{dt}=0$. This results in a set of algebraic relationships given by the solutions of
    $\mathbf{S}_{\mathbf{y}}\mathbf{v}\left(\mathbf{y},\mathbf{z}\right)=0$ (23)
    where $\mathbf{S}_{\mathbf{y}}$ is the submatrix containing only the reactions involving $\mathbf{y}$. Within the SODAs framework, the goal is to identify the DAE with algebraic equations $\mathbf{G}=\mathbf{S}_{\mathbf{y}}\mathbf{v}=0$ and dynamic equations $\mathbf{F}=\frac{d\mathbf{z}}{dt}=\mathbf{S}_{\mathbf{z}}\mathbf{v}$, where $\mathbf{S}_{\mathbf{z}}$ contains only the reactions involving $\mathbf{z}$. Substitution of algebraic relationships via the implicit definition of $\mathbf{y}$ in terms of $\mathbf{z}$ allows for the expression of the system entirely in terms of $\mathbf{z}$. With $\mathbf{y}=\mathbf{g}(\mathbf{z})$, CRN dynamics are given by
    $\frac{d\mathbf{x}}{dt}=\frac{d\mathbf{z}}{dt}=\mathbf{S}_{\mathbf{z}}\mathbf{v}(\mathbf{g}(\mathbf{z}),\mathbf{z}).$ (24)
    As all terms in the non-reduced model are polynomial, the reduced model takes the rational form in (10). This form was the target of discovery by implicit-SINDy [26] and PI-SINDy [27]. An advantage of discovering the DAE form rather than the related rational equations is that we discover which variables are in quasi-steady state rather than assuming that all measured variables are inherently dynamic. Note that this means that SODAs requires measurements of both algebraic and differential variables, which is unlikely for realistically-sized biological or chemical networks. Future work to lessen the burden of unobserved states is discussed in Section 5. In realistic examples, the method in its current form may be best-suited instead for model reduction during the simulation of large CRNs, which is an ongoing area of study [60, 61]. For definitions of $\mathbf{S}_{\mathbf{z}}$ and $\mathbf{v}$ in all tested CRNs, see supplementary Section 2.1.
@@
    We start with the discovery of the equations of CRN1, the single irreversible Michaelis-Menten reaction. Within the seemingly simple dynamics, there exist three distinct algebraic relationships, two induced by mass conservation and one by separating the state variables $\mathbf{x}$ as $\mathbf{x}=[\mathbf{y},\mathbf{z}]^{\top}$, resulting in a system of the form in (23)-(24). Of the two mass conservation relationships, one is of the form (15) (i.e., conservation of enzyme) and the other relates $[A]$ and $[B]$ with a constant $M^{\text{tot}}$ representing total non-enzyme mass (i.e., conservation of substrate). The single algebraic relationship emerging from the quasi-steady-state approximation takes the implicit form of
    $[AE_{1}]=\frac{\left(E_{1}^{\text{tot}}-[AE_{1}]\right)[A]}{K_{M1}+[A]}$ (25)
    where $K_{MN}$ is the Michaelis constant for reaction $N$ derived from the set $\{k\}$ of the reaction’s rate constants from (21). Expanding the above implies an algebraic relationship between terms $[AE_{1}]$, $[AE_{1}][A]$, and $[A]$ as in (19).
    To discover such relationships from our simulated time series, we first constructed a candidate library using a monomial basis of degree $p$ in states $\mathbf{x}$. Such a basis allows for the discovery of two of the three algebraic relationships when multiple simulations of the CRN are used. Because $M^{tot}$ varied between simulations, identification of the substrate conservation relationship was not possible in our setup, as a single learned $M^{tot}$ cannot fit all simulations. This was not the case for the enzyme conservation relationship, as $E_{\text{tot}}$ was held constant between simulations. Inclusion of non-zero initial conditions in $\Theta$ would allow for the discovery of substrate conservation, a close analog to the inclusion of suspected bifurcation parameters in $\Theta$ in the original formulation of SINDy [10] and following work. We chose here to exclude initial conditions from $\Theta$ because when any two (or more) initial conditions are held constant across simulations, they will be in perfect correlation, confounding one another and limiting the identifiability and interpretability of any resulting algebraic relationships. Because substrate mass conservation can be checked by simply observing whether the system is open or closed, we deemed it reasonable that this third algebraic relationship does not benefit from sparse regression for its discovery. If this is not a reasonable assumption, for example in a large or particularly complex system, note that simulations or experiments may be designed alternatively such that initial conditions in $\Theta$ do not have these colinearities (see Section 5).
@@
    We performed discovery for CRN1 using degree $p=2,3,4$ libraries with a range of 0-15% noise to assess the data requirements for successful recovery of algebraic relationships with SODAs (Figure 3b). In the context of CRNs, this gives a reference for the quality and resolution of data needed to determine whether mass is conserved among a subset of states or if there is timescale separation in certain reactions sufficient for quasi-steady-state approximations to hold. For each application of SODAs, data was preprocessed using Savitzky-Golay filters [48] to smooth noisy derivatives. We used the results of the Pareto front approach referenced in Section 3 to determine that the library should be refined two times. Since we were interested not simply in the minimum data to achieve rediscovery of the correct algebraic equations, but the minimum data to ensure rediscovery at a given resolution and any resolution finer, we defined the SODAs data requirement to be the number of uniformly spaced data points per initial condition above which algebraic relationship recovery was ensured up to a tested maximum of 1200 data points per initial condition. Note that this does not prevent time series with lower data resolution than those indicated in Figure 3b from being used for successful model discovery with SODAs. In low-resolution cases, the impact of noise on the regression problem is heightened, meaning that correct discovery may be possible but is dependent on the particular realization of measurement error (see supplementary Section 2.2). Subsequent discovery of the dynamic equations returned the correct structure and parameter sufficiently similar to those used for simulation (see supplementary Section 2.4 for parameter fitting results at varied noise).
@@
    All algebraic relationships were correctly recovered with up to $15\%$ noise for all degrees of $p$ tested. Even for large magnitudes of noise, SODAs required a lower resolution and far fewer sampled initial conditions for recovery than implicit-SINDY [26] or PI-SINDY [27]. The amount of data required for recovery across library degrees scaled similarly for low noise, but the introduction of higher-order terms in the $p=3$ and $p=4$ libraries lead to a sharp increase in the data requirement as noise becomes significant. Due to corruption of the signal and noise amplification, these higher-order terms induce spurious algebraic relationships with correlations similar to those of true relationships for low time-series resolutions. For both $p=3$ and $p=4$, as the noise was increased beyond $12\%$ and $6\%$ respectively, recovery failed to be robust. This suggests that there are certain noise realizations for which $\leq 1200$ data points are sufficient for rediscovery, but it should not be expected.
    Testing recovery in the more complex CRN2 and CRN3 examples demonstrates how SODAs responds to a larger state space, richer dynamics, and, as a consequence, more complex algebraic relationships. In the tested CRNs of higher dimension, the same type of relationships exist as in CRN1: each enzyme is conserved, and each reaction equilibrates in a manner that makes a quasi-steady-state approximation reasonable. However, because of coupling between the reactions, these algebraic relationships can appear with an increased complexity and number of terms, where it is not immediately obvious the exact physical property from which they result. Take, for example, the quasi-steady-state approximations made in CRN2 (Figure 3a). Quasi-steady-state approximations for both the $B\to C$ and $B\to D$ reactions mediated by enzymes $E_{2}$ and $E_{3}$ respectively, give the implicit relationships
    $[BE_{2}]$ $=\frac{\left(E_{2}^{\text{tot}}-[BE_{2}]\right)[B]}{K_{M2}+[B]}$ $[BE_{3}]$ $=\frac{\left(E_{3}^{\text{tot}}-[BE_{3}]\right)[B]}{K_{M3}+[B]}$ (26)
    in the same form as (25) once we assume that $[E_{2}]$ and $[E_{3}]$ have been removed from the library by a previous iteration of SODAs due to mass conservation. Note that the choice to remove enzymatic terms instead of their complexes is a standard choice to derive the Michaelis-Menten dynamics [37, 38], but either the enzyme or the complex could be removed according to the SODA algorithm, as they are of equal order. Isolating $[B]$ in each expression in (26), we equate to find
    $2(K_{M3}-K_{M2})[BE_{2}][BE_{3}]=K_{M3}E_{2}^{\text{tot}}[BE_{3}]-K_{M2}E_{3}^{\text{tot}}[BE_{2}]$ (27)
    which suggests that the quasi-steady-state approximation relationships should ultimately appear in the candidate library as an algebraic relationship in terms of $[BE_{2}][BE_{3}]$, $[BE_{3}]$, and $[BE_{2}]$. An analogous derivation exists for CRN3.

### v2

@@ 3 Methods: Sparse Optimization for Differential-Algebraic Systems
For the identified differential variables, any existing methods for ODE discovery (e.g., SINDy) can be used to find the dynamics equations in the DAE system (Figure 1c). This is the stage where noisy derivatives affect recovery, and there are various approaches to mitigate noise, including weak formulation approaches [Messenger and Bortz [46, 44], Wentz and Doostan [80]] and smoothing functions [Sun et al. [73], Bromba and Ziegler [15]]. We have used Savitzky-Golay filters [Bromba and Ziegler [15]] in our examples. For numerical experiments, we have used LASSO followed by sequential thresholding to solve the sparse optimization. Finally, we obtain a set of differential and algebraic equations of the form (2)–(3) (Figure 1d).
Input: $\mathbf{X}$, $K$, $\Theta$, $\alpha^{a}$, $\alpha^{d}$, sparse_fit, score_function, $\epsilon$
@@
$k=0$  %set iteration index
condition_number_improved = True   %flag for condition number improvement.
@@
[…]
@@
 $m=\arg\max\limits_{l}(S^{k}_{l})$  %find the best relationship based on the maximum score.
@@
 $\theta^{(k)}_{\text{high}}=\texttt{select\_term}(g_{k}(\mathbf{x}))$  %select the term in $g_{k}(\mathbf{x})$ based on complexity rank.
 $\Gamma_{k}=\texttt{multiples}(\theta^{(k)}_{\text{high}},\Theta_{k-1})$   %identify all multiples of $\theta^{(k)}$ in $\Theta_{k-1}$.
@@
[…]
K=k1K=k-1
$\Theta_{*}=\Theta_{k-1}$%setting final library to the one .
1exStep 3: Dynamical System Finder
@@
$\mathbf{F}(\dot{\mathbf{y}},\ddot{\mathbf{y}},\mathbf{x},\ldots)=0\xleftarrow{}\texttt{discover\_ODEs}(\mathbf{y},\Theta_{*},\alpha^{d})$   %discovering ODEs using refined library.
1ex
@@
Details on the inputs and defaults can be found in Supplementary-section S1.
@@ 4 Results
We demonstrate the capability of SODAs to discover DAEs and identify reduced coordinate systems through three distinct examples. In Example 1, SODAs rediscovers the underlying DAEs of CRNs with varying complexity, demonstrating the identification of conservation laws and quasi-steady states. In Example 2, SODAs rediscovers DAEs used in power-grid modeling for three benchmarks (IEEE-4, IEEE-9, and IEEE-39) demonstrating recovery of network topology when the library contains trigonometric functions. Example 3 applies SODAs to pixel data from real-time experiments and animated video corresponding to three different cases of nonlinear pendulum: a singular pendulum with negligible damping, a heavily damped pendulum, and a chaotic double pendulum. This example differs from the first two in that it focuses on identifying the algebraic constraints and, therefore, a reduced coordinate system, which is not readily apparent from pixel data. For the single-pendulum example, we also demonstrate recovery of a dynamic model in the new coordinate.
In Examples 1 and 2, where the time series were generated synthetically, we asses data requirements, robustness to noise, and sensitivity to different types of perturbations are studied. For a simulated time-series states, $\mathbf{X}_{i}$ , Gaussian noise is sampled and added to the time series as
@@
where $noise\%$ is the percentage of noise, $\sigma_{i}=\text{std}(\mathbf{X}_{i})$, and $\mathcal{N}(0,1)$ is the standard normal random variable. In Example 3, the pixel data was extracted from video footage, which inherently introduced noise into the dataset. In Examples 1 and 3, we explore the impact of polynomial library size on data requirements. For both the algebraic finder and the ODE discovery step, we used LASSO followed by sequential thresholding as the default sparse solver in all our numerical experiments, unless specified otherwise. The values of the hyperparameters representing sparsity enforcement penalties, $\alpha^{a}$ and $\alpha^{d}$, for each numerical experiment where applicable, are provided in the Supplementary-section S6.
@@ 4.1 Example 1: Chemical Reaction Networks
We demonstrate the ability of SODAs to accurately recover governing equations from simulated time series of known CRNs, and study data requirements as system complexity and extrinsic noise increase. In applications in chemistry and biology, the temporal evolution of each chemical species in a CRN can be modeled using mass-action kinetics [Érdi and Tóth [22], Alon [1]]. This naturally leads to models consisting of polynomial terms of the concentrations of chemical species. However, timescale separations in the equilibration rates of different species motivate quasi-steady-state approximations [Bowen et al. [12], Segel and Slemrod [66]] which simplify mathematical analysis by reducing the state space dimension [Goeke et al. [26]].
An advantage of discovering the DAE form rather than the related rational equations is that we discover which variables are in quasi-steady state rather than assuming that all measured variables are inherently dynamic. For known systems, models can be simplified through conservation analysis, partitioned into sub-networks, or non-dimensionalized. Discovering the algebraic systems from data could automate the process of model reduction in chemical and biological networks when the systems are unknown. Beyond quasi-steady-state approximations, other optimization and sensitivity analysis can motivate simplification [Snowden et al. [70]]. Previous work used neural network-based model discovery as a framework for CRN reduction [Bhattacharya et al. [10], Benner et al. [7]]. SODAs demonstrates identifies likely conservation-of-mass relationships and timescale separation using regression-based sparse-model selection, resulting in an interpretable reduction when model structure is unknown. Note that SODAs requires measurements of both algebraic and differential variables, which is unlikely for realistically-sized biological or chemical networks. Future work to lessen the burden of unobserved states is discussed in Section 5. In realistic examples, the method in its current form may be best-suited instead for model reduction during the simulation of large CRNs, which is an ongoing area of study [Yang et al. [82], Wen et al. [79]].
@@
In this example, we use time series generated from three CRNs of varying complexity consisting of one, two, and four enzyme-mediated reactions (Figure 3a). The four-reaction system describes the simplified propanediol utilization pathway native to Salmonella enterica [Mills et al. [49], Kennedy et al. [35]]. CRN1 is the single irreversible reaction (4- 7.) CRN2 and CRN3 follow a similar mass action kinetic form (full models, parameters, and initial conditions in S3). CRN1 has two algebraic relationships, one representing conservation of enzyme (6) and one from a quasi-steady state approximation (7). Conservation equations are common in modeling CRNs, representing the notion that enzymes are not created or degraded on short time-scales, and that the system is closed [22]. CRN2 has four algebraic and three dynamic variables. CRN3 has nine algebraic and five dynamic variables. Because of the reversibility of the $B]\Leftrightarrow[C]$ reaction in CRN3, the algebraic relationships have increased complexity. For example, for $E_{2}]$, enzyme conservation results in $E_{2}]+[BE_{2}]+[CE_{2}]-E_{2}^{tot}=0$ and the quasi-steady-state approximations $k_{4}[B][E_{2}]-(k_{5}+k_{8})[BE_{2}]+k_{9}[CE_{2}]\approx 0$. Irreversible reactions in CRN2 and CRN3 follow very similar forms to those introduced in CRN1.
During the simulation of the systems to produce time-series to test SODA’s recovery, we assume the algebraic relationships hold exactly using the systems in reduced Michaelis-Menten form (i.e. 9 and 10, and S3(c)). Each system was simulated from five different initial conditions, varying the initial concentration of the first substrate in each network, so that the range of dynamic behavior was similar in all instances explored. As SODAs has no knowledge of the difference between algebraic relationships arising from conservation or quasi-steady-state these equations can appear as any linear combination of the equations (See S3(b)). In the dynamic finder step, we expect to find equations for the dynamics of the substrates. For example in CRN1 this is $A$, and the product $B$, represented by (4) and (5) (See S3(a) for CRN2 and CRN3 systems).
Figure 3: Example 1: Application to chemical reaction networks. (a) Nested structure of CRN1 and CRN2 within CRN3 and examples of simulated time-series with $5\%$ noise. Note that only CRN3 allows reversibility of $C\to BE_{2}$ as denoted by the green arrow. (b) Scaling of the sampling rate required for recovery of the correct algebraic equations in CRN1 across library degree $p$ and cardinality $|\Theta|$. Non-robust recovery is denoted when fewer than $80\%$ of random seeds succeeded in recovering the correct algebraic relationships. Error bars represent standard error across 10 random noise seeds. (c) Data requirement for recovery of the correct algebraic equations for all CRNs (cardinalities noted) with fixed library degree $p=2$. Each system instance used five initial conditions which are listed in Supplementary-section S3(c).
@@
We tested SODAs recovery of the expected algebraic relationships from the simulated time series of CRN1, using libraries of polynomials with degree $p=2,3,4$ and a range of 0-15% noise (Figure 3b). The high end of this noise range is similar to realistic noise in metabolite measurements [Bartha [6], Meyer et al. [47]]. Data was preprocessed using Savitzky-Golay filters [Bromba and Ziegler [15]] to smooth noisy derivatives. Since we were interested in the minimum data to ensure rediscovery for each library and noise condition, we defined the SODAs data requirement to be the smallest number of uniformly spaced data points per initial condition above which the algebraic relationships were recovered for 10 noise instances. The time series generated from five initial conditions were sampled at varied frequencies up to a maximum of 1200 data points per initial condition (see Supplementary-section S3(c) for visualizations of time series). At lower sampling resolution than indicated in Figure 3b, recovery may be possible but is dependent on the particular realization of measurement error (see Supplementary-section S3(d)). Subsequent discovery of the dynamic equations returned the correct structure and parameter sufficiently similar to the ground truth (see Supplementary-section S3(f) for examples of parameter fitting results at varied noise).
@@
The algebraic relationships were correctly recovered with up to $15\%$ noise for the degrees of $p$ tested for at least some noise instances. For $p=2$ and up to 11% noise for $p=3$, recovery was robust across all noise instances. The amount of data required for recovery across library degrees scaled similarly for low noise, but the introduction of higher-order terms, especially in the $p=4$ library, led to a sharp increase in the data requirement as noise becomes significant. Due to corruption of the signal and noise amplification, these higher-order terms likely induce spurious algebraic relationships with correlations similar to those of true relationships. For $p=4$, as the noise was increased beyond $6\%$, recovery failed to be robust. There are certain noise realizations for which $\leq 1200$ data points are sufficient for rediscovery, but rediscovery is not reliable. SODAs required a lower resolution of data and only five initial conditions for recovery compared to implicit-SINDy [Mangan et al. [42]] or SINDy-PI [Kaheman et al. [33]], which demonstrated recovery using time-series from over 3,000 initial conditions. We also attempted to apply PI-SINDy which is the most robust algorithm among implicit SINDy methods, to discover the models using data where SODAs was successful, but were unable to recover the correct dynamics even when using the maximum amount of data considered in this study (Figure 3b: 5 initial conditions and 1200 time-series points per initial condition) with as low as $1\%$ noise. The method was applied using the PySINDy implementation of PI-SINDy (details in Supplementary-section S3(g)). However, SODAs requires knowledge of the time-series of complexes and enzymes (ie. $AE_{1}]$ and $E_{1}]$), while SINDy-PI requires only the timeseries of the substrates (ie. $[A]$), so the comparison is not direct and challenging to interpret.